### v2

# Rapid and Safe Trajectory Planning over Diverse Scenes through Diffusion Composition

Abstract
Safe trajectory planning in complex environments must balance stringent collision avoidance with real-time efficiency, which is a long-standing challenge in robotics. In this work, we present a diffusion-based trajectory planning framework that is both rapid and safe. First, we introduce a scene-agnostic, MPC-based data generation pipeline that efficiently produces large volumes of kinematically feasible trajectories. Building on this dataset, our integrated diffusion planner maps raw onboard sensor inputs directly to kinematically feasible trajectories, enabling efficient inference while maintaining strong collision avoidance. To generalize to diverse, previously unseen scenarios, we compose diffusion models at test time, enabling safe behavior without additional training. We further propose a lightweight, rule-based safety filter that, from the candidate set, selects the trajectory meeting safety and kinematic-feasibility requirements. Across seen and unseen settings, the proposed method delivers real-time-capable inference with high safety and stability. Experiments on an F1TENTH vehicle demonstrate practicality on real hardware. Project page: https://rstp-comp-diffuser.github.io/.
I Introduction
Traditional trajectory planners for complex environments typically comprise two phases: initial path generation and subsequent trajectory refinement via optimization. While search- and sampling-based methods can produce initial paths [1], their runtime grows rapidly with scene complexity [2], and the resulting paths are often not kinematically feasible [3]. The subsequent optimization step can enforce feasibility and safety, but adds computational burden. For time-critical applications such as planning for autonomous vehicles, the key challenge is to jointly achieve low computational cost, kinematic feasibility, and reliable collision avoidance.
Imitation learning offers an effective alternative to classical two-stage planners by directly learning from optimization-generated trajectories and modeling the distribution of near-optimal solutions [4], thereby avoiding repetitive global search and post hoc optimization. Among imitation learning methods, the diffusion model is widely used due to its ease of training [5], robust generation sampling process [6], and natural support for multi-modal inputs [7]. However, the diffusion model suffers from slow trajectory generation due to its iterative denoising process, which can lead to failure in real-world deployment. In addition, such imitation learning methods heavily rely on the training data – datasets of suboptimal kinematics or limited diversity can easily compromise performance.
To this end, we propose a diffusion-based Rapid and Safe Trajectory Planning (RSTP) framework that includes an efficient MPC-based dataset generation method, an integrated diffusion planning pipeline that maps raw sensor input to executable trajectories, and a real-world compositional novel scene generalization strategy.
The proposed MPC-based dataset generation method uses differential equations to describe system dynamics, thus eliminating extensive reinforcement learning agent training and enabling plug-and-play use across different scenes. As MPC enforces kinematic constraints, the resulting trajectories are kinematically feasible and well-suited as expert demonstrations. Furthermore, to enable responsive diffusion planning in real-time control scenarios, we introduce a fast trajectory‐synthesis framework that seamlessly integrates real‐time perception with diffusion‐based generative models. By fusing low-level vehicle sensory inputs using Simultaneous Localization and Mapping (SLAM) into compact vehicle states, our approach dramatically accelerates the iterative denoising process while fully preserving the diffusion model’s capacity to produce precise, high‐fidelity trajectories.
In addition, traditional global trajectory planning often requires substantial adjustments to handle diverse scenarios. Compared to traditional planning methods, the proposed RSTP framework can seamlessly adapt to diverse scenarios without additional tuning or demonstrations by leveraging test-time diffusion composition. Specifically, Fig. LABEL:fig:vis_abstract (a) and (b) depict two diffusion models trained separately on a static and a dynamic scene. In an unseen composite scenario, for example, where both static obstacles from (a) and a dynamic obstacle from (b) are presented, our method can still generate safe trajectories by composing these two diffusion models. As shown in Fig. LABEL:fig:vis_abstract (c), this test-time diffusion composition yields adaptive behaviors, such as accelerating to bypass obstacles or decelerating to avoid obstacles with no extra model training.
To harden deployment, we propose a lightweight safety filter to further enhance planning safety in real-world applications. We validate the proposed pipeline on the popular F1TENTH platform [8], which is an efficient testbed for planning methods that integrates perception, planning, and control in a unified hardware system. Experiments on a custom F1TENTH vehicle confirm the superior real-world performance of our framework.
In summary, the proposed RSTP framework based on the diffusion model has the following contributions:
  1. 1.
    Feasibility-Aware, Data-Driven Trajectory Planning: We introduce an end-to-end training pipeline that includes an energy-based diffusion trajectory planner and a data generation method based on MPC. We demonstrate that such a framework is able to learn the distribution of collision-free, kinematically feasible trajectories, yielding diffusion samples that are safe and directly trackable without post hoc optimization.
  2. 2.
    Rapid and Safe Diffusion Planning from Raw Sensor Inputs: By operating in a low-dimensional vehicle-state extracted from perception, RSTP achieves 0.21s mean planning time on commodity hardware. Furthermore, a rule-based safety filter selects an optimal trajectory from a candidate batch, ensuring safety.
  3. 3.
    Real-time Diffusion Composition for Flexible Planning in Diverse Scenes: We formulate a composition mechanism that enables generalization to planning in unseen scenarios and demonstrate its effectiveness in time-critical decision-making scenarios with noisy real-world sensory inputs.
  4. 4.
    Validation on a Real-World Scaled F1TENTH Vehicle Platform: The proposed method is validated on an F1TENTH vehicle using server-based inference with simple onboard pure pursuit tracking. The vehicle follows the planned trajectory accurately and without collisions, demonstrating the kinematic feasibility and real-world applicability of our approach.
This paper is organized as follows: Section II introduces the use of diffusion models to represent the planning problem. Section III explains the method for dataset generation. Section IV provides a detailed description of the composition of diffusion models for diverse scenes and the safety filter. Section V presents numerical simulations and experiments based on the F1TENTH platform. Section VI concludes the paper.
II Preliminaries
In this section, the method of path planning using potential fields [9] and its drawbacks are first introduced. Then, the approach of using Energy-based Models (EBMs) for potential field modeling is discussed in  [10]. Finally, the relationship between diffusion models and EBMs is explored.
II-A Path Planning using Potential Field
Collision-free path planning can be achieved using the potential-based approach [10]. In this method, a potential function $\mathcal{U}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is used to assign a potential value to a path $\bm{\tau}$, based on predefined attractive and repulsive potentials. Here, $n$ represents the state dimension and $L$ is the path horizon. Given an initial path $\bm{\tau}$, the collision-free path can be derived by applying gradient descent to the potential function as follows: $\bm{\tau}^{k}_{1:L}=\bm{\tau}^{k-1}_{1:L}-\lambda\nabla_{\bm{\tau}_{1:L}^{k}}\mathcal{U}(\bm{\tau}_{1:L}^{k}),$ where $\bm{\tau}^{k}_{1:L}$ is the path at step $k$, and $\bm{\tau}^{k-1}_{1:L}$ is the path at the previous step, and $\lambda$ is the step size for an iteration. However, this method leads to corner cases, making it unsuitable for complex scenarios.
II-B Energy-based Models
Energy-based Models are a class of generative models [11]. They can serve as surrogate models for the potential function, represented by $\mathcal{U}(\tau)\sim E_{\theta}(\bm{\tau})$, addressing the limitations of conventional potential-based methods. Given a path variable $\bm{\tau}$, the probability density $p_{\theta}(\bm{\tau})$ of a collision-free path $\bm{\tau}$ is defined as: $p_{\theta}(\bm{\tau})\propto e^{-E_{\theta}(\bm{\tau})}$, where $E_{\theta}(\bm{\tau}):\mathbb{R}^{n\times L}\to\mathbb{R}$ is an energy-based model, typically implemented using a U-Net architecture. Thus, the path planning process can be modified as follows: $\bm{\tau}_{1:L}^{t-1}=\bm{\tau}_{1:L}^{t}-\lambda\nabla_{\bm{\tau}_{1:L}^{t}}E_{\theta}(\bm{\tau}_{1:L}^{t},t,\mathcal{C})+\eta,$ where $\eta\sim\mathcal{N}(\mathbf{0},{\sigma}_{t}^{2}\bm{I})$, with $\lambda$ and $\sigma_{t}$ being specific scaling values. $t$ is the denoising step. And $\mathcal{C}=\{\mathbf{s}_{s},\mathbf{s}_{g},\mathcal{O}\}$ is the set of constraints, which includes the start state and goal state constraints denoted as $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$, as well as the obstacle pose constraints $\mathcal{O}$. In this paper, $\tau=[\mathbf{s}_{1},\cdots,\mathbf{s}_{L}]_{4\times L}$ denotes a discrete planned path. Specifically, $\mathbf{s}=[x,\,y,\,q_{z},\,q_{w}]^{\top}$ with $x,\,y]^{\top}$ representing the Cartesian position and $q_{z},\,q_{w}]^{\top}$ the components of the quaternion encoding the yaw angle.
II-C Diffusion Models
Diffusion models $\epsilon_{\theta}$ effectively estimate the energy-based model’s gradient, denote as $\nabla_{\bm{\tau}^{t}_{1:L}}E_{\theta}(\bm{\tau}^{t}_{1:L},\mathcal{C})\propto\epsilon_{\theta}(\bm{\tau}^{t}_{1:L},t,\mathcal{C})$. The loss function for a single diffusion model training is:
$\mathcal{L}_{\mathrm{MSE}}=\mathbb{E}_{p(\bm{\tau}),\,\epsilon}\|\epsilon-\epsilon_{\theta}(\bm{\tau}^{t}_{1:L},t,\mathcal{C})\|^{2}$
where $\epsilon\sim\mathcal{N}(\mathbf{0},\bm{I})$ represents Gaussian noise. $\bm{\tau}_{1:L}^{t}$ is the generated path at the denoising step $t$.
The forward noise addition process is given by: $q(\bm{\tau}_{1:L}^{t}\mid\bm{\tau}_{1:L}^{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\bm{\tau}_{1:L}^{t},\,\beta_{t}\bm{I}),$ where $\beta_{t}$ is the variance schedule. After a sufficient number of noising steps $T$, the final distribution of $q(\bm{\tau}^{\top}_{1:L})$ converges to: $q(\bm{\tau}^{\top}_{1:L})\approx\mathcal{N}(\mathbf{0},\bm{I}).$ On the other hand, the reverse denoising process can be expressed as: $p_{\theta}(\bm{\tau}_{1:L}^{t-1}\mid\bm{\tau}_{1:L}^{t}):=\mathcal{N}(\mu_{\theta}(\bm{\tau}_{1:L}^{t},t),\,\tilde{\beta_{t}}^{2}\bm{I}),$ where $\mu_{\theta}(\bm{\tau}_{1:L}^{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{\tau}_{1:L}^{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\bm{\tau}_{1:L}^{t},t)\right),$ and $\alpha_{t},\bar{\alpha}_{t},\tilde{\beta_{t}}$ are functions of $\beta_{t}$.
III MPC-based Data Collection
Fig. 3 overviews the RSTP pipeline, which proceeds in three modules. The first module, a comprehensive data collection procedure, is described in detail in this section.
For collision avoidance in the static scene, global collision-free paths are generated using the iterative collision avoidance (ItCA) method [3]. The ItCA path planning accounts for the smoothness of gear shifting points (GSPs), ensuring the kinematic feasibility of planned paths. Then, an MPC planner is employed for path tracking to produce the final global trajectory. The entire trajectory enriches the static dataset, implicitly encoding velocity through the density of trajectory points within a fixed interval $\Delta t$. Implicit learning of velocity enables more complex trajectories, exhibiting behaviors such as decelerating to yield or accelerating to bypass.
The vehicle model $f(\cdot)$ used for MPC is the kinematic model employed in [2]. The state is represented as $\bm{\zeta}=[x,\,y,\,\varphi,\,v]^{\top}$ and the control input as $\mathbf{u}=[\delta,\,a]^{\top}$. $\delta$ is the steering angle, and $a$ represents acceleration. $v$ is the velocity and $\varphi$ represents the vehicle heading.
Figure 2: Diffusion Composition. Trajectory planning using the energy model as a surrogate model for the potential field.
For global collision-free ItCA path planning, starting states $\mathbf{s}_{s}$ are arbitrarily sampled within the feasible region. The goal states $\mathbf{s}_{g}$ are fixed for all candidate points. The velocity of both $\mathbf{s}_{s}$ and $\mathbf{s}_{g}$ is 0. The reference path pose sequence generated by the ItCA method, $\mathbf{p}_{s_{k}}^{\text{ref}}=[x_{s_{k}}^{\text{ref}},\,y_{s_{k}}^{\text{ref}},\,\varphi_{s_{k}}^{\text{ref}}]^{\top}$, is parameterized by the arc length $s_{k}$.
In dynamic scenes, a penalty term based on the distance between the ego vehicle and dynamic obstacles is added to the cost function: $J_{\textrm{dynamic}}=\alpha\sum_{i=1}^{N_{\mathrm{o}}}({\|\mathbf{p}_{\text{cur}}-\mathbf{o}_{i}\|_{2}})^{-1}$, where $N_{o}$ is the number of dynamic obstacles and $\mathbf{o}_{i}$ is the pose of the $i$-th obstacle. $\alpha$ is the coefficient. $\mathbf{p}_{\text{cur}}$ is the current vehicle pose. It is emphasized that dynamic obstacle avoidance does not consider static obstacles, which may result in collisions with them in the planned trajectory.
The improved Euler method is used to discretize the kinematic model $f(\bm{\zeta})$, where $\bm{\zeta}_{0}$ represents the current state of the vehicle. Then the optimization problem for the MPC planner across diverse scenes is expressed as follows:
$\min_{\bm{\zeta},\mathbf{u}}$ $\gamma\cdot J_{\textrm{dynamic}}+(1-\gamma)\cdot\Big(\sum_{k=1}^{N_{p}}\overbrace{\left(\|\mathbf{A}\cdot\bm{\zeta}_{k}-\mathbf{p}_{s_{k}}^{\text{ref}}\|^{2}_{Q_{1}}\right)}^{J_{\textrm{static}}}$ (1)
$+\sum_{k=1}^{N_{p}-1}\underbrace{\|\Delta\mathbf{u}_{k}\|^{2}_{R_{1}}+\|\mathbf{A}\cdot\bm{\zeta}_{N_{p}}-\mathbf{p}_{s_{N_{p}}}^{\text{ref}}\|^{2}_{Q_{2}}+\|\mathbf{u}_{k}\|^{2}_{R_{2}}\Big)}_{J_{\textrm{static}}}$
s.t. $\bm{\zeta}_{k}=\bm{\zeta}_{k-1}+T_{s}\cdot f\left(\bm{\zeta}_{k}+\frac{T_{s}}{2}f\left(\bm{\zeta}_{k},\mathbf{u}_{k}\right),\mathbf{u}_{k}\right),$
$\bm{\zeta}_{0}=\bm{\zeta}_{\textrm{cur}},\quad\bm{\zeta}_{\textrm{min}}\leq\lvert\bm{\zeta}_{k}\rvert\leq\bm{\zeta}_{\textrm{max}},\quad\mathbf{u}_{\textrm{min}}\leq\lvert\mathbf{u}_{k}\rvert\leq\mathbf{u}_{\textrm{max}},$
where $\mathbf{A}=[1,\,1,\,1,\,0]^{\top}$, and $N_{p}$ denotes the prediction horizon. $\bm{\zeta}_{\textrm{cur}}$ is the current vehicle state. In dynamic scenes, $\gamma=1$; otherwise, $\gamma=0$. Following the concept of projected velocity in [12, 13], the value of $s_{k}$ is determined by a fixed projection velocity $v_{\text{ref}}$, such that $s_{k}=s_{k-1}+\Delta t\cdot v_{\text{ref}}$, with $s_{0}=s_{\text{cur}}$. $\bm{\zeta}_{\text{max}}$ and $\bm{\zeta}_{\text{min}}$ denote the upper and lower bounds of the state constraints, while $\mathbf{u}_{\text{max}}$ and $\mathbf{u}_{\text{min}}$ are the upper and lower limits of the control inputs, respectively. $T_{s}$ represents the prediction interval. $Q_{1}$, $Q_{2}$, $R_{1}$, and $R_{2}$ are coefficient matrices.
To ensure a consistent trajectory horizon $L$ during dataset generation, horizons shorter than $L$ are padded by duplicating the last value, whereas longer trajectories are truncated.
Ultimately, the collision-free demonstration trajectories for the ii‑th given start and end states is represented as $\bm{\tau}_{i}^{\text{dem}}=[x_{i},y_{i},q_{i,z},q_{i,w}]^{\top}_{4\times L}.$ In dynamic scenarios, the recorded obstacle trajectory is represented as $\mathcal{O}^{\text{single}}_{i}=[x_{i,\text{obst}},y_{i,\text{obst}}]_{2\times L}^{\top}$. The final trajectory dataset is represented as $D=\{{\bm{\tau}_{i}^{\text{dem}},\mathcal{O}^{\text{single}}}\}_{i=1}^{N_{D}},$ where $N_{D}$ is the number of candidate starting points. A key aspect of generating the dataset is ensuring that the time interval $\Delta t$ remains constant, which is especially critical for collision avoidance.
IV The Proposed Rapid and Safe Trajectory Planning Method
This subsection provides a detailed description of the second and third modules in Fig. 3. The second module achieves generalization to unseen scenes by composing the individually trained diffusion models. The third module applies a safety filter to guarantee the safety of the final planned trajectory.
Figure 3: The Overall Framework of the Proposed Rapid and Safe Trajectory Planning (RSTP) Method. Offline Dataset Generation (Left): the ItCA [3] and MPC-based methods provide kinematically feasible trajectory datasets for training. Diffusion Composition (Middle): individual diffusion models can be flexibly composed to tackle novel scenarios not covered in the training data. Online Inference and Control (Right): The ego vehicle performs real-time inference from its current pose, and the safety filter selects the optimal trajectory per cycle for tracking.
IV-A Diffusion Models Composition
The probability density of an explicit collision-free trajectory $\bm{\tau}_{1:L}$, conditioned on separate static or dynamic obstacle constraints $\mathcal{O}_{i}$, is denoted by $p^{\text{traj}}_{\theta_{i}}(\bm{\tau}_{1:L}\mid\mathcal{O}_{i})$. For unseen scenarios, specifically when dynamic obstacles appear in a static environment, the probability density of the optimal trajectory is represented as $p^{\text{traj}}_{\theta_{\text{unseen}}}(\bm{\tau}_{1:L}\mid\mathcal{O}_{\text{all}})$, where $\mathcal{O}_{\text{all}}=\{\mathcal{O}_{\text{static}},\mathcal{O}_{\text{dyn}}\}$ denotes the union of the static and dynamic obstacle positions. In both static and dynamic scenarios, constraints on the start and goal states are enforced and, for clarity, are omitted from the formula.
The trained models corresponding to the dynamic and static scenarios can be composed to form the final model $p^{\text{traj}}_{\theta_{\text{unseen}}}$. This process is denoted as:
$p^{\text{traj}}_{\theta_{\text{unseen}}}(\bm{\tau}_{1:L}$ $\mid\mathcal{O}_{\text{all}})\propto p^{\text{traj}}_{\theta_{\text{compose}}}(\bm{\tau}_{1:L},\mathcal{O}_{\text{static}},\mathcal{O}_{\text{dyn}})$ (2)
$\propto p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L})\prod_{i=1}^{N_{\text{all}}}\frac{p^{\text{traj}}_{\theta_{i}}(\bm{\tau}_{1:L}\mid\mathcal{O}_{i})}{p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L})}$
where $N_{\text{all}}$ is the number of models to be composed and $\mathcal{O}_{i}\in\mathcal{O}_{\text{all}}$. The obstacles $\mathcal{O}_{i}$ are encoded using a Transformer architecture as adapted from [10]. The probability density of a collision-free trajectory is referred to as the unconditional probability density $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$. This distribution is directly learned from the dataset without explicit obstacle constraints. Each $p^{\text{traj}}_{\theta}$ corresponds to a diffusion model. And the $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$ and conditional diffusion model $p^{\text{traj}}_{\theta}(\bm{\tau}_{1:L}^{t}\mid\mathcal{O}_{i})$ can be obtained by masking the obstacle-encoder input with $\mathbf{K}=[1,0]^{\top}\otimes\mathcal{O}_{i}$ during inference, allowing the same model to output both conditional and unconditional results.
The training of the diffusion model can be viewed as the gradient of an energy model, where satisfies $\nabla_{\bm{\tau}^{t}_{1:L}}\log p^{\text{traj}}_{\theta}(\bm{\tau}^{t}_{1:L}\mid\mathcal{O}_{i})\propto\epsilon^{\text{traj}}_{\theta}(\bm{\tau}^{t}_{1:L},t,\mathcal{O}_{i})$. Thus, Eq. (2) can be rewritten as:
$\epsilon^{\text{traj}}_{\theta_{\text{compose}}}$ $(\bm{\tau}_{1:L}^{t},t,\{\mathcal{O}_{i}\}_{i=1}^{N_{\text{all}}})=\epsilon^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}^{t}_{1:L},t,\emptyset)$ (3)
$+\sum_{i=1}^{N_{\text{all}}}\nu_{i}\cdot\left(\epsilon^{\text{traj}}_{\theta_{i}}(\bm{\tau}^{t}_{1:L},t,\mathcal{O}_{i})-\epsilon^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}^{t}_{1:L},t,\emptyset)\right)$
where $\nu_{i}\in\bm{\nu}_{1\times N_{\text{all}}}$ is the composition parameter and $t$ is the denoising time step. $\epsilon^{\text{traj}}_{\theta_{i}}$ is the trained diffusion model based on different obstacle distributions.
Within a reasonable range, the larger the value of $\nu_{i}$, the more the final composed trajectory will resemble the trajectory generated by the model $\epsilon^{\text{traj}}_{\theta_{i}}$. Therefore, this allows for diffusion composition that enables the planning of safe global trajectories in unseen scenarios with extra obstacles.
IV-B Safety Filter
In test-time planning, denoising from random Gaussian noise introduces variations in the planned trajectories, which can potentially cause instability and even collisions. To ensure safety, a batch of trajectories is generated, and the highest-quality one is selected as the final trajectory. The overall process is in Algorithm 1.
The batch of all trajectories is denoted as $\mathcal{B}^{\text{test}}=\{\mathbf{q}_{i},\bm{\varphi}_{i}\}_{i=1}^{N_{\text{filter}}}$, where $\mathbf{q}_{i}=[x,\,y]^{\top}_{2\times L}$ denotes the point sequence of the $i$-th trajectory, and $N_{\text{filter}}$ is the number of trajectories. The $\mathcal{B}^{\text{test}}_{i}=\{\mathbf{q}_{i},\bm{\varphi}_{i}\}$ can be derived from the planned trajectory $\bm{\tau}_{i}=[x_{i},y_{i},q_{i,z},q_{i,w}]^{\top}_{4\times L}$.
First, the kinematic feasibility is considered. The distance between adjacent points in $\mathbf{q}_{i}$ is given by: $\mathbf{d}_{i}={\left(\Delta\mathbf{q}_{i,[1,:]}^{2}+\Delta\mathbf{q}_{i,[2,:]}^{2}\right)^{\frac{1}{2}}}$, where $\Delta\mathbf{q}_{i}=\mathbf{q}_{i,[:,2:L]}-\mathbf{q}_{i,[:,1:L-1]}$. Then the trajectory length of $\mathbf{q}_{i}$ can be given by $l_{i}=\sum_{j=1}^{L-1}d_{i,j}$. It is used as a metric because minimizing the trajectory length helps to avoid redundant GSPs. The velocity $\mathbf{v}_{i}$ and acceleration $\mathbf{a}_{i}$ are then calculated as: $\mathbf{v}_{i}=\frac{\mathbf{d}_{i}}{\Delta t}\,\text{and}\,\mathbf{a}_{i}={(\mathbf{v}_{i+1}-\mathbf{v}_{i})}/{\Delta t}$. Then the yaw rate is given by: $\mathbf{r}={\left(\bm{\varphi}_{i,[2:L]}-\bm{\varphi}_{i,[1:L-1]}\right)}/{\Delta t}$. Finally, the corresponding steering angle is: $\bm{\delta}_{i}=\arctan\left(\frac{l_{w}\cdot\mathbf{r}_{i}}{\mathbf{v}_{i}}\right)$, where $l_{w}$ is the wheelbase of the vehicle.
To guarantee numerical stability, we normalize the cost terms so their scales remain uniform. After obtaining $\mathbf{L}=\{l_{i}\}_{i=1}^{N_{\text{filter}}}$, $\mathbf{a}=\{\mathbf{a}_{i}\}_{i=1}^{N_{\text{filter}}}$, and $\bm{\delta}=\{\bm{\delta}_{i}\}_{i=1}^{N_{\text{filter}}}$ for the trajectory batch $\mathcal{B}^{\text{test}}$, the corresponding values are normalized. The normalization process for $\mathbf{L}$ is given by $\frac{\mathbf{L}-\min(\mathbf{L})}{\max(\mathbf{L})-\min(\mathbf{L})}.$ For acceleration batch $\mathbf{a}$, the normalization is: $\frac{\mathbf{\|a\|}^{2}-\min(\mathbf{\|a\|}^{2})}{\max(\mathbf{\|a\|}^{2})-\min(\mathbf{\|a\|}^{2})}.$ The same normalization process is applied to $\bm{\delta}$. Then the final normalized values are denoted as $\tilde{\mathbf{L}}$, $\tilde{\mathbf{a}}$, and $\tilde{\bm{\delta}}$, respectively.
To ensure safety, a minimum distance, $\rho_{i}$, is enforced between the trajectory $\mathbf{q}_{i}$ and obstacles. This distance acts as a penalty metric for trajectories that come too close to obstacles. The corresponding cost function is:
$J_{i,\text{safe}}=\frac{1}{d_{i}+1}\cdot\mathbb{I}(\rho_{i}>0)+V_{\text{inf}}\cdot\mathbb{I}(\rho_{i}=0)$
where $\mathbb{I}(\cdot)$ is the indicator function, which is 1 when the condition is satisfied and 0 otherwise. A value of $\rho_{i}=0$ indicates a collision with obstacles, in which case $J_{i,\text{safe}}$ is assigned a large penalty value $V_{\text{inf}}$ to discard the trajectory effectively. Finally, the overall cost function as the safety filter is as follows:
$J_{i}=\omega_{1}\tilde{\mathbf{L}}_{i}+\omega_{2}\|\tilde{\mathbf{a}}_{i}\|+\omega_{3}\|\tilde{\bm{\delta}}_{i}\|+\omega_{4}J_{i,\text{safe}},\forall\,i\in[1,N_{\text{filter}}]$ (4)
where $\omega_{1}$ to $\omega_{4}$ are the corresponding weights. Each of the cost terms in (4) is normalized, allowing their corresponding weights to be easily tuned in different scenarios. $J_{\text{best}}=\min\{J_{i}\}_{i=1}^{N_{\text{filter}}}$. The smaller the value of $J_{i}$, the higher the quality of the trajectory. The final computed value of $J_{\text{best}}$ corresponds to a single trajectory $\mathcal{B}$, which demonstrates superior safety, smoothness, and efficiency among a batch of candidate trajectories.
To prevent abrupt changes in the yaw angle along the selected trajectory $\mathcal{B}$, a post hoc check is conducted on the variation $\Delta\bm{\varphi}=\bm{\varphi}_{[2:L]}-\bm{\varphi}_{[1:L-1]}$ between adjacent trajectory points. If any $\Delta\bm{\varphi}$ exceeds $w_{max}\cdot\Delta t$, where $w_{\max}$ denotes the vehicle’s maximum angular velocity, the corresponding metric $\phi(\mathcal{B})$ is assigned a significant penalty value $V_{\text{inf}}$; otherwise, it is set to 0. This mechanism is designed to filter out trajectories that violate motion constraints, thereby ensuring kinematic feasibility, which is denoted as follows:
$\phi(\mathcal{B})=0,~\text{if }\Delta\bm{\varphi}_{[1:L]}<w_{max}\cdot\Delta t,\;\text{otherwise}\;V_{\text{inf}}$ (5)
For execution safety, if the number of replanning attempts exceeds $N_{\text{retry}}$ without finding a feasible solution, the planner returns an empty result and the vehicle refrains from acting.
Algorithm 1 The Safety Filter for Diffusion Planning.
0: diffusion model: $\epsilon^{\text{traj}}_{\theta}$,  candidate batch : $N_{\text{filter}}$, number of retries: $N_{\text{retry}}$, constraints: $\mathbf{s}_{s}$, $\mathbf{s}_{g}$,  $\mathcal{O}_{\text{all}}$.
[…]
2:while $k\leq N_{\text{retry}}$ do
3:  $\mathcal{B}^{\text{test}}\leftarrow$ generate $N_{\text{filter}}$ inference trajectories simultaneously using $\epsilon^{\text{traj}}_{\theta}$ constrained on $\mathbf{s}_{s}$, $\mathbf{s}_{g}$ and $\mathcal{O}_{\text{all}}$;
4:  $J_{\text{best}}$,  $\mathcal{B}$$\leftarrow$ find the minimum cost $J_{\text{best}}$ in $\mathcal{B}^{\text{test}}$ and its corresponding trajectory; % Apply safety filter and obtain the optimal trajectory in the batch.
5:  if $J_{\text{best}}<V_{\text{inf}}$ and $\phi(\mathcal{B})<V_{\text{inf}}$ then % Motion limits.
6:  $\quad\tau^{\star}\leftarrow{\mathcal{B}}$;
7:   Return $\tau^{\star}$; % Find the optimal trajectory.
8:  end if
9:  $k\leftarrow k+1$; % Replanning.
10:end while
11: Return $\emptyset$;
11: The finally planned trajectory $\tau^{\star}$ with safety.
V EXPERIMENTAL RESULTS AND DISCUSSION
The proposed RSTP method is validated through both simulation and real-world tests. In simulation, individual model and their compositions are evaluated using safety, computational efficiency, and kinematic feasibility metrics. The safety filter’s effectiveness is also confirmed in simulation. Trajectories generated by the composed models are then tested on the F1TENTH vehicle platform and show strong safety and kinematic feasibility under real-world uncertainties.
V-A Environments and Baselines
Details of the experimental setup are provided in Appendix A. The diffusion model adopts the architecture and hyperparameters of [10], and its composition formula follows (3). In all experiments, the weights of the terms in Eq.(4) are set equal. The following scenes and experimental setups are established for simulation and real-world testing:
  • 1)
    Static Scene ($SS$): Includes only static obstacles. The diffusion model trained on the static scene is referred to as the Static Model (StM).
  • 2)
    Composed Scene 1 ($CS_{1}$): Features the same static obstacles as SS, along with a single moving obstacle vehicle. This vehicle is initialized at $x_{1}=3.5,y_{1}=1.2,q_{z}=1.0,q_{w}=0.00]$ and moves in a straight line at a constant speed of $0.4\text{m/s}$ for $7.1\,\text{s}$, as shown in Fig. 6(a). The diffusion model trained only on this dynamic obstacle trajectory in $CS_{1}$ is denoted as Dynamic Model 1 (DyM1).
  • 3)
    Composed Scene 2 ($CS_{2}$): Features the same static obstacles as SS, along with a single moving obstacle vehicle. This vehicle is initialized at $x_{2}=3.5,y_{2}=2.0,q_{z}=0.891,q_{w}=-0.454]$ and moves in a straight line at a constant speed of $0.4\text{m/s}$ for $5.3\,\text{s}$, as shown in Fig. 6(b). The diffusion model trained only on this dynamic obstacle trajectory in $CS_{2}$ is denoted as Dynamic Model 2 (DyM2).
We compare our RSTP method with multiple existing approaches. Baseline methods for comparison include $\text{A}^{\star}$[14], $\text{RRT}^{\star}$[15], and $\text{FMT}^{\star}$[16], all of which disregard vehicle kinematics. Additionally, kinematics-aware methods, Kino RRT[17] and $\text{Hybrid A}^{\star}$[1] are included for comparison. We denote the baseline method $\text{Hybrid A}^{\star}$ as BLM. The optimization-based methods evaluated are Rapid Iterative Trajectory Planning (RITP) [3] and Optimization-Based Collision Avoidance (OBCA) [2], both of which are initialized using paths generated by $\text{Hybrid A}^{\star}$. To assess kinematic feasibility, the MPC tracker described in Section IV is employed as the tracking controller. The mean tracking error (M.TE) serves as the metric for evaluating kinematic feasibility. Computational efficiency is evaluated by measuring the computation time (C.T). For the diffusion model, the C.T is the time to generate the final collision-free trajectory, including replanning. The percentage reduction in C.T compared to the BLM is denoted as $\text{M.RP}=\frac{\text{C.T}_{\text{BLM}}-\text{C.T}}{\text{C.T}_{\text{BLM}}}$. Similarly, the percentage reduction in M.TE is represented by $\text{M.RP}=\frac{\text{M.TE}_{\text{BLM}}-\text{M.TE}}{\text{M.TE}_{\text{BLM}}}$. The configuration with the safety filter disabled is indicated as RSTP-NSF.
TABLE I: Simulation Results and Evaluation Metrics for the RSTP Method: Demonstrating Time Efficiency and Safety.
Methods4 F.Rate1 C.Rate2 C.T M.TE Max Min Mean M.RP Std Max Min Mean M.RP Std Proposed RSTP 0.57% 8.80% 0.522 0.137 0.210 98.88% 3.16e-4 12.909 0.321 1.424 47.23% 1.647 RSTP-NSF 4.50% 11.87% 0.515 0.144 0.195 98.96% 1.65e-4 37.933 0.284 1.495 44.62% 4.299 Heuristics $\text{Hybrid A}^{\star}$ 24.88% 27.75% 248.964 0.004 18.747 - $\uparrow$ 6.929 0.459 2.699 - 0.882 Kino RRT 14.64% 58.18% 39.362 0.019 4.267 77.24% 31.601 39.686 0.476 3.232 -19.77% 16.876 $\text{A}^{\star}$ 0.86% 67.37% 5.337 0.138 2.032 89.16% 0.924 53.129 1.085 3.903 -44.62% 8.772 $\text{RRT}^{\star}$ 1.53% 59.43% 13.950 0.160 3.177 83.05% 3.029 23.561 0.951 4.290 -58.96% 12.235 $\text{FMT}^{\star}$ 0.77% 74.07% 2.766 0.615 1.450 92.27% 0.124 10.194 0.284 5.188 -92.23% 4.237 Optimization3 RITP 0.76% 7.87% 0.351 0.008 0.059 99.69% 0.002 6.413 0.254 0.792 70.67% 0.377 OBCA 10.41% 13.20% 1.395 0.378 0.508 97.29% 0.005 5.877 0.178 0.560 79.27% 0.111
  • 1 The Failure Rate (F.Rate) is defined as $\frac{N_{f}}{N_{D}^{\text{all}}}$, where $N_{f}$ denotes the number of planning failures. The planning failure is defined as the planner’s failure to generate a collision-free path or trajectory within a specified number of iterations and time.
  • 2 The Collision Rate (C.Rate) is defined as $\frac{N_{f}+N_{c}}{N_{D}^{\text{all}}}$, where $N_{c}$ represents the number of trajectories that collides during tracking.
  • 3 For optimization-based methods, $\text{Hybrid A}^{\star}$ is used for warm-start, incurring an additional mean computation time of 18.747 s.
  • 4 Trajectories are planned in the Static Scene $SS$. The RSTP and RSTP-NSF are based on the StM.
V-B Evaluate the Performance of RSTP Method in Simulation
We first verify the safety and efficiency of RSTP, then validate the safety filter, and finally assess how the composition of diffusion models generalizes to new scenes. Experimental results are summarized in Table I. The sensitivity of weight settings for diffusion composition is discussed in Appendix -C.
V-B1 On the Safety Performance of our RSTP method
The results summarized in Table I show that the proposed RSTP achieves a Failure Rate (F.Rate) of only 0.57% and a low Collision Rate (C.Rate) of 8.8% during tracking. By implicitly learning the vehicle kinematics constraints, RSTP significantly reduces collisions during tracking, outperforming kinematics-aware methods such as Hybrid $\text{A}^{\star}$ and Kino RRT, which suffer from higher failure rates of 24.88% and 14.64% due to the limitations of discrete search and sampling in long-horizon planning. Optimization-based methods, RITP and OBCA, explicitly enforce kinematics and achieve comparable collision rates of 7.87% and 13.2%, compared to RSTP. However, they require well-defined warm-start paths, which increase computation time. More analysis on the kinematic feasibility of planned trajectories is available at Appendix -B. In summary, RSTP delivers notable safety and efficiency gains, benefits to the diffusion model’s generative ability, and stability.
V-B2 On the Time Efficiency of our RSTP method
As shown in Table I, the proposed RSTP achieves a mean C.T of only 0.21s across all starting poses in the evaluation dataset $D_{A}$, representing a 98.88% improvement of computational efficiency compared to BLM. Additionally, it demonstrates strong stability, with a standard deviation of only 3.16e-4, outperforming all comparison methods. This time efficiency stems from state-based denoising, which significantly reduces the noise dimension. The time stability benefits from the strong stability of diffusion models in generating trajectories. Optimization-based methods, such as RITP and OBCA, depend on Hybrid $\text{A}^{\star}$ for initialization. This dependency results in an additional mean computational cost of 18.747 s. In contrast, the RSTP method avoids this extra warm-start overhead. In conclusion, state-based diffusion inference facilitates rapid trajectory planning, enabling the vehicle to respond quickly to environmental conditions with strong stability.
V-B3 On the Effectiveness of Safety Filter
Compared to RSTP, the F.Rate and C.Rate of RSTP-NSF increase to 4.50% and 11.87%, respectively, emphasizing the necessity of utilizing the safety filter. Additionally, the mean C.T for RSTP-NSF is 0.195 s, comparable to the 0.21 s measured with the safety filter enabled. This indicates minimal computational overhead introduced, thus validating its suitability for time efficiency. Furthermore, the standard deviation of M.TE increases from 1.647 in RSTP to 4.299 in RSTP-NSF, indicating a decrease in weak kinematic feasibility.
((a))
((b))
Figure 4: Diffusion Model Can Ensure Kinematic Constraints. The RSTP method can plan safe trajectories involving large curvature 4(a) and gear shifting points (GSP) 4(b).
Figure 5: The Composed Model Flexibly Avoids Obstacles in Unseen Scenarios by Adjusting Velocity. Comparison of F1TENTH’s velocity under different diffusion models in $CS_{1}$. The composed model is able to decelerate to avoid static obstacles, whereas the dynamic model (DyM1) fails to avoid the additional static obstacles.
V-C Evaluate Compositional Performance in Real-world
The real-world experiments are conducted using a self-built F1TENTH vehicle. For mapping and localization, the vehicle utilizes the Cartographer ROS framework [18]. Communication between the server and onboard computer is managed via the Message Queuing Telemetry Transport (MQTT) protocol. The F1TENTH vehicle uses the pure pursuit controller proposed in [19] to track trajectories. The RSTP method can plan sufficiently safe trajectories in the $SS$ as shown in Fig. 4.
The effectiveness of the model composition for safety is first validated through trajectory planning in the real world. Fig. 5 clearly shows the difference in velocities between the composed model and the individual models. The trajectory generated by the DyM1 alone results in a collision with a static obstacle. However, when composing the StM with the DyM1, the StM steers the planned trajectory of DyM1, effectively avoiding collisions with static obstacles. This generalization capability is clearly evident from the decision to decelerate, as marked in Fig. 5. Fig. 6 presents the collision-free trajectories generated by the composed models in scenes $CS_{1}$ and $CS_{2}$. In both scenarios, the composed models successfully produce trajectories that avoid collisions. These experimental results demonstrate that the model composition enhances the safety of the RSTP method for real-world applications.
To further evaluate the safety of the proposed RSTP method, three distinct starting points are selected within obstacle-free regions of the map in both scenes $CS_{1}$ and $CS_{2}$. The safety metric used is the minimum distance between the vehicle vertices and both static and dynamic obstacles, with results summarized in Table II. Even with localization uncertainty, the minimum obstacle distance recorded is 0.032 m, confirming adequate safety margins. Moreover, fewer than 10% of trajectory points fall within the defined danger threshold, indicating that only a minor segment of the planned trajectory poses a potential collision risk. Thus, the composed model proves to be sufficiently safe in the context of real-world uncertainty.
((a))
((b))
Figure 6: Safe Planning with Composition. Schematic diagram of trajectories planned by different models in composed scenes. Green denotes the ego trajectory; red denotes the dynamic-obstacle trajectory.
TABLE II: The Vehicle–Obstacle Distances in Real-world Execution when Following the Trajectory Generated by the Composed Model.
Scenes Start points Danger1 Max Min Mean Std $CS_{1}$ [0.86, 0.03] 6.60% 0.591 0.032 0.283 0.020 [3.03, -0.07] 8.54% 0.559 0.038 0.260 0.020 [2.73, 0.32] 3.77% 0.532 0.054 0.288 0.016 $CS_{2}$ [0.82, 0.07] 4.55% 0.590 0.051 0.278 0.015 [2.76, 0.47] 0.00% 0.593 0.086 0.304 0.019 [2.79, -0.11] 7.41% 0.550 0.036 0.275 0.018
  • 1 The Danger metric is defined as ${N_{\text{Danger}}}/{N_{all}}$, where $N_{\text{Danger}}$ is the number of points with a minimum distance to other obstacles below 0.08 m, and $N_{all}$ is the total number of points.
VI Conclusion
This paper introduces the RSTP framework, which combines a diffusion model-based planner with a test-time safety filter. The key contributions include a real-time, fast diffusion planner integrated with onboard perception; the MPC-based dataset generation; a model composition strategy that enables safe generalization to unseen scenarios with high stability; and extensive validation on a real-world F1TENTH vehicle. Experimental results demonstrate the effectiveness of RSTP and highlight the impact of model composition weights on performance. Future work will extend RSTP to racing overtaking scenarios and incorporate constraints-in-the-loop.
This appendix is structured as follows: Section -A details the experimental setup and hardware configurations; Section -B examines the kinematic feasibility of trajectories generated by the RSTP method; and Section -C evaluates the performance of the model composition within simulation.
-A Experiment
To evaluate the RSTP, training and evaluation datasets are constructed from a real-world $6\,\text{m}\times 6\,\text{m}$ map. From the map’s free navigable area, a comprehensive evaluation dataset $D_{A}$ containing $N_{D}^{\text{all}}=1045$ uniformly sampled initial poses is created. To specifically assess the compositional model, a subset $D_{C}$ with $N_{D}^{\text{comp}}=79$ initial poses is selected from $D_{A}$. Target poses remain fixed. The trajectory horizon $L$ is set to 128 with 100 diffusion steps during training. Denoising employs the DDIM method with 8 steps, and the number of candidate trajectories $N_{\text{filter}}$ for the safety filter is set to 8. Training and inference are performed using an RTX 4060 GPU and a 4.9 GHz Intel(R) i7-12700 CPU with 48GB RAM.
-B On the Kinematic Feasibility of our RSTP method
The proposed RSTP method also demonstrates excellent kinematic feasibility, as presented in Table I. Specifically, the M.TE of RSTP shows a substantial reduction of 47.23% compared to BLM. This significant improvement indicates that the diffusion model can effectively and implicitly learn vehicle kinematics from the dataset. In contrast, methods such as $\text{A}^{\star}$, $\text{RRT}^{\star}$, and $\text{FMT}^{\star}$, which neglect vehicle kinematic constraints, lead to considerably larger tracking errors. Optimization-based methods, such as RITP and OBCA, achieve notable reductions in M.TE of 70.67% and 79.27%, respectively, yielding trajectories with strong kinematic feasibility. However, this improved accuracy comes at a high computational cost. In comparison, RSTP generates trajectories with strong kinematic feasibility while demanding lower computational resources.
-C Effectiveness of Model Composition and Sensitivity Analysis of Compositional Weights
The effectiveness of diffusion models composition for generalization is validated in scenes $CS_{1}$ (Fig. 6(a)) and $CS_{2}$ (Fig. 6(b)), employing four distinct compositional weight configurations, as summarized in Table III. As indicated in the table, the composed model can safely generalize to unseen scenes when appropriate compositional weights are selected, achieving an F.Rate of 0%. However, increasing the weight assigned to conditional models results in a rise in F.Rate across both test scenes. These results demonstrate that the generalization capability is sensitive to compositional weights, and maintaining appropriate weights is critical for minimizing failures.
TABLE III: Simulation Evaluation using Models Composition in Out-of-Distribution Scenes of the RSTP Method.
Scenes cond uncond-dynamic1 cond uncond-static1 F.Rate C.T F.Rate C.T Mean Std Mean Std $CS_{1}$ 0.8 0.00% 0.643 0.100 0.3 0.00% 0.618 0.002 2.5 13.92% 1.039 1.157 2.5 53.16% 1.730 3.382 5.0 34.18% 1.800 3.930 5.0 43.04% 2.181 3.013 8.0 51.90% 5.630 30.508 8.0 32.91% 5.179 38.695 $CS_{2}$ 0.8 0.00% 0.653 0.011 0.2 0.00% 0.622 0.019 2.5 43.04% 1.446 2.679 2.5 68.35% 2.199 4.812 5.0 40.51% 3.554 11.316 5.0 64.56% 4.498 19.689 8.0 69.62% 13.163 91.913 8.0 53.16% 6.360 31.140
  • 1 The compositional weight of all unconditional diffusion models is 5.